Two asteroids are separated by a distance of 42.0 m. If one asteroid has a mass of 2.22 x 109 kg and the other asteroid has a mass of  9.87 x 108 kg, what is gravitational force between them?



1
Expert's answer
2020-12-25T14:07:51-0500

The gravitational force between two asteroids is given by the Newton's law:


"F = G\dfrac{m_1m_2}{r^2}"

where "m_1 = 2.22\times 10^9 kg", "m_2 = 9.87\times 10^8kg" are masses of the asteroids, "r = 42m" is the distance between the asteroids, and "G = 6.67\times 10^{-11}m^3kg^{-1}c^{-2}". Combining it all together, obtain:


"F = 6.67\times 10^{-11}\cdot \dfrac{2.22\times 10^9\cdot 9.87\times 10^8}{42^2} \approx 7.73\times 10^4N"

Answer. "7.73\times 10^4N".